Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2019-08-20

Observers: Morgan Fraser, Seán Brennan
Support Team: Nada Ihanec, Marius Gromadzki
TAT: Steve Williams, Claudia Gutierrez, Thomas Wevers

Calibrations: 
EFOSC2:
biases
dome flat and arcs
UBVRi skyflats (R and i have long exposures - will repeat tomorrow)


Weather: 
Start: Humidity 10%, wind speed 5 m/s, seeing: 0.9, clear
Middle:
End: 


Comments:

Instrument - EFOSC2

UT            Target                        Setup                                    Comments
23:06        EG274                        Gr13+11+16                          Skipped image analysis so that we could get on science targets sooner.
23:22        SN2019gfm                Gr13                                      Followup, barely visible on acquisition image - S/N~1 in 2700s spectrum
00:20        SN2019fdr                Gr16                                        Followup
01:16                                                                                       Telescope error
01.28        SN2019fdr                Gr11                                       Followup
02:21        SN2019cqc               Gr13                                    First spectrum aborted as on wrong target
03:20                                                                                       Image analysis
03:27        SN2019hcc              griz                                        Folllowup
04:22        AT2019nln               Gr13                                     Classification - low S/N and ambiguous               
04:55        AT2019noh             Gr13                                      Classification
05:18        AT2019nyn              Gr13                                     Classification
05:29        AT2019nyk              Gr13                                      Classification
05:30        SN2019mrz             Gr13                                       Folllowup
06:17        SN2019mry             Gr11+Gr16                          Followup
08:04         AT2019nog                                                           Classification, target not visible on 60s acquisition
08:13        AT2019mic               Gr13                                         Classification
08:38       AT2019nwf             Gr13                                      Classification
08:49                                                                                    Switch to SOFI
08:57       SN2019mry            J                                           Followup - sequence interrupted by telescope error at exposure 10 of 12.
09:12                                                                                    Telescope error - rotator stopped
09:15       SN2019mry            HK                                       Recovered from error, resumed with H band
09:24                                                                                     Telescope error - rotator stopped
09:30       SN2019mry            HK                                       Recovered from error, restarted H band
10.26       SN2019esa           JHK

Comments